OpenStudy (anonymous):

Lisa can make a costume twice as fast as Justin can. If they both work together, they can make it in 8 hours. How long will it take Mary to mike the costume if he works alone?

OpenStudy (queelius):

First, I assume you mean "How long will it take Justin to make the costume if he works alone?" So, here's how I thought about it. Suppose a costume consists of n units of work. So, together, Lisa (L units/hour) and Justin (J units/hour) can make n units in 8 hours: L*8 + J*8 = n We are given that L = 2J. Thus: 3J*8 = n => J = n/24 So, L = n/12. Lisa by herself, then, would take L*h = n => n/12*h=n => 12/h=1 => h=12. So, 12 hours.

OpenStudy (anonymous):

ohh. i understand now. thanks for your help! @queelius

OpenStudy (queelius):

You're welcome.